OpenStudy (anonymous):

integral of (x^3/4) - (1/(x^3/4)) dx

sam (.sam.):

\[\int\limits_{}^{}x^{3/4}-\frac{1}{x^{3/4}}dx\] \[\int\limits_{}^{}x^{3/4}-{x^{-3/4}}dx\] \[\frac{x^{7/4}}{7/4}-\frac{x^{1/4}}{1/4}\]

sam (.sam.):

\[\frac{4 x^{7/4}}{7}-4 \sqrt[4]{x}+c\] \[\frac{4}{7} \sqrt[4]{x} \left(x^{3/2}-7\right)+c\]